# STONE: Self-supervised Tonality Estimator

### v1

@@ Abstract
Although deep neural networks can estimate the key of a musical piece, their supervision incurs a massive annotation effort. Against this shortcoming, we present STONE, the first self-supervised tonality estimator. The architecture behind STONE, named ChromaNet, is a convnet with octave equivalence which outputs a key signature profile (KSP) of 12 structured logits. First, we train ChromaNet to regress artificial pitch transpositions between any two unlabeled musical excerpts from the same audio track, as measured as cross-power spectral density (CPSD) within the circle of fifths (CoF). We observe that this self-supervised pretext task leads KSP to correlate with tonal key signature. Based on this observation, we extend STONE to output a structured KSP of 24 logits, and introduce supervision so as to disambiguate major versus minor keys sharing the same key signature. Applying different amounts of supervision yields semi-supervised and fully supervised tonality estimators: i.e., Semi-TONEs and Sup-TONEs. We evaluate these estimators on FMAK, a new dataset of 5489 real-world musical recordings with expert annotation of 24 major and minor keys. We find that Semi-TONE matches the classification accuracy of Sup-TONE with reduced supervision and outperforms it with equal supervision.
@@ 1 Introduction
By adding an extension to STONE, we present Semi-TONE, a semi-supervised model that is tailored for 24-way key estimation. Semi-TONE performs on par with a supervised counterpart (Sup-TONE) while reducing dependency on annotated data by 90% [^1] https://github.com/Yuexuan-Kong/STONE (will be ready soon).
@@
A new large dataset of 5489 real-world music recordings, collected from the Free Music Archive (FMA) and annotated by an expert for 24 major and minor keys, is available for free download [^2] FMAK dataset: https://zenodo.org/records/10719860.
@@ 2.1 Equivariant self-supervised learning in music
Historically, SSL techniques for MIR learn general-purpose embeddings which need a fine-tuning stage to address downstream tasks. Yet, In recent years, a new family of methods has emerged: i.e., learning how to solve a specific task by representing the transformations which underlie its factors of variability. Within the realm of music signal processing, several studies have applied equivariant self-supervised learning for pitch and tempo estimation[10, 11, 12]. In particular, PESTO is a monophonic pitch estimator trained by learning the pitch shift of the same sample[9]. However, its extension to multipitch tracking is an open problem [13].
@@ 3.1 Artificial pitch transpositions of the CQT
Given a CQT matrix $\bm{x}$ and some integer $c\leq 15$, we reduce the number of rows from 99 down to 84 by trimming the $c$ lowest-frequency bins and $(15-c)$ highest-frequency bins. This is tantamount by a pitch transposition by $c$ semitones. We denote the result by $T_{c}\bm{x}$ where $T_{c}\bm{x}[p,t]=\bm{x}[p-c,t]$ for each frequency $p<84$ and time $t$.
@@ 3.2 ChromaNet: a convnet with octave equivalence
We compose the convnet $f_{\bm{\theta}}$ with a non-trainable operator $g$ whose role is to guarantee octave equivalence. We roll the log-frequency axis into a spiral which makes a full turn at every octave, thus aligning coefficients in $f_{\bm{\theta}}(T_{c}\bm{x})$ of the form $(p\pm Qj)$ for integer $j$. The operator $g$ sums these coefficients across octaves $j$ for each pitch class $q$ applies a softmax transformation. Composing $f_{\bm{\theta}}$ and $g$ yields
@@ 3.3 DFT over key signature profiles
We interpret $\bm{y}_{\bm{\theta}}$ as a learned key signature profile (KSP). With $Q=12$, its discrete Fourier transform (DFT) is
@@
where $\omega$ is an integer between 00 and $11$ that is co-prime to 12 for a full circular distribution over all 12 pitches. With $\omega=7$, a circular shift of $\bm{y}_{\bm{\theta}}$ by seven chromas corresponds to a multiplication of $\widehat{\bm{y}}_{\bm{\theta}}[\omega]$ by $e^{2\pi\mathrm{i}49/12}=e^{2\pi\mathrm{i}/12}$. Hence, the phase of the complex number $\widehat{\bm{y}}_{\bm{\theta}}[\omega]$ denotes a key modulation in the circle of fifths (CoF). Alternatively, $\omega=1$ would correspond to a circle of semitones. Our paper evaluates both settings but only describes the CoF setting ($\omega=7$) for the sake of conciseness.
@@ 5.3 Self-supervised mode estimation
We now introduce a noncontrastive loss for self-supervised mode estimation. To this aim, we posit that mode is not only constant throughout the musical piece, but also remains invariant by pitch transposition. Therefore, going back to the notations from Section 4: $T_{c}\bm{x}_{\mathrm{A}}$, $T_{c}\bm{x}_{\mathrm{B}}$, and $T_{c+k}\bm{x}_{\mathrm{B}}$ should elicit the same value of the mode estimator $\mu_{\bm{\theta}}$.
@@ 6.3 Results on key and mode classification
Training Semi-TONE on the full GSMK dataset yields a MIREX score of 68.7%; i.e., roughly on par with the supervised SOTA (69.2%). Figure 4 shows the confusion matrix of calibrated STONE and Semi-TONE on FMAK. Although our methods do not outperform the SOTA on key estimation, it brings insights into a novel framework that does not require high supervision for training. Moreover, we note that self-supervision remains beneficial even when the full GSMK dataset is available for training. Therefore, a promising avenue of research is to scale up the dataset of unlabeled recordings (see Section 4.1), thus widening the gap between Semi-TONE and Sup-TONE on FMAK.
@@ 7 Conclusion
STONE demonstrates that its key signature profile learned in a fully unsupervised manner could correlate with key signatures and perform well for key signature estimation without any supervision. Moreover, STONE exhibits its transferability to key estimation and shows that semi-TONE could reach comparable results as sup-TONE with 90% less labeled data. The primary limitation of STONE lies in training the model in a fully self-supervised way to distinguish major keys from minor keys. As of now, the 24-STONE are not yet comparable to models with minimal supervision. Future work will study how STONE can adapt to other pitch-related MIR tasks such as chord detection, and the generalization across culturally diverse musical corpora.

### v2

@@ Abstract
Although deep neural networks can estimate the key of a musical piece, their supervision incurs a massive annotation effort. Against this shortcoming, we present STONE, the first self-supervised tonality estimator. The architecture behind STONE, named ChromaNet, is a convnet with octave equivalence which outputs a “key signature profile” (KSP) of 12 structured logits. First, we train ChromaNet to regress artificial pitch transpositions between any two unlabeled musical excerpts from the same audio track, as measured as cross-power spectral density (CPSD) within the circle of fifths (CoF). We observe that this self-supervised pretext task leads KSP to correlate with tonal key signature. Based on this observation, we extend STONE to output a structured KSP of 24 logits, and introduce supervision so as to disambiguate major versus minor keys sharing the same key signature. Applying different amounts of supervision yields semi-supervised and fully supervised tonality estimators: i.e., Semi-TONEs and Sup-TONEs. We evaluate these estimators on FMAK, a new dataset of 5489 real-world musical recordings with expert annotation of 24 major and minor keys. We find that Semi-TONE matches the classification accuracy of Sup-TONE with reduced supervision and outperforms it with equal supervision.
@@ 1 Introduction
We extend STONE into Semi-TONE, a semi-supervised model that is tailored for 24-way key estimation. Semi-TONE performs on par with a supervised counterpart (Sup-TONE) while reducing dependency on annotated data by 90% [^1] Companion website: https://github.com/deezer/stone.
@@
A new large dataset of 5489 real-world music recordings, collected from the Free Music Archive (FMA) and annotated by an expert for 24 major and minor keys, is available for free download [^2] FMAK dataset: https://zenodo.org/records/12759100.
@@ 2.1 Equivariant self-supervised learning in music
Equivariant SSL learns task-specific embeddings by representing the transformations which underlie its factors of variability: e.g., variations in pitch or tempo [10, 11, 12]. In particular, PESTO is a monophonic pitch estimator trained by learning the pitch shift of the same sample[9]. However, its extension to multipitch tracking is an open problem [13].
@@ 3.1 Artificial pitch transpositions of the CQT
Given a CQT matrix $\bm{x}$ and a integer $c\leq 15$, we reduce the number of rows from 99 down to 84 (7 octaves) by trimming the $c$ lowest-frequency bins and $(15-c)$ highest-frequency bins. This is tantamount by a pitch transposition by $c$ semitones [9]. We denote the result by $T_{c}\bm{x}$ where $T_{c}\bm{x}[p,t]=\bm{x}[p-c,t]$ for each frequency $p<84$ and time $t$.
@@ 3.2 ChromaNet: a convnet with octave equivalence
We compose the convnet $f_{\bm{\theta}}$ with a non-trainable operator $g$ whose role is to guarantee octave equivalence. We roll the log-frequency axis into a spiral which makes a full turn at every octave, thus aligning coefficients in $f_{\bm{\theta}}(T_{c}\bm{x})$ of the form $(p\pm Qj)$ for integer $j$. This averaging process is a common operation in literature[26]. The operator $g$ sums these coefficients across octaves $j$ for each pitch class $q$ and applies a softmax transformation. Composing $f_{\bm{\theta}}$ and $g$ yields
@@ 3.3 DFT over key signature profiles
We propose the term “key signature profile” for the learned vector $\bm{y}_{\bm{\theta}}$. With $Q=12$, its discrete Fourier transform (DFT) is
@@
where $\omega$ is an integer between 00 and $11$ that is coprime to 12 for a full circular distribution over all 12 pitches. With $\omega=7$, a circular shift of $\bm{y}_{\bm{\theta}}$ by seven chromas corresponds to a multiplication of $\widehat{\bm{y}}_{\bm{\theta}}[\omega]$ by $e^{2\pi\mathrm{i}49/12}=e^{2\pi\mathrm{i}/12}$. Hence, the phase of the complex number $\widehat{\bm{y}}_{\bm{\theta}}[\omega]$ denotes a key modulation in the circle of fifths (CoF). Alternatively, $\omega=1$ would correspond to a circle of semitones. Our paper evaluates both settings but only describes the CoF setting ($\omega=7$) for the sake of conciseness.
@@ 5.3 Self-supervised mode estimation
We now introduce a loss for self-supervised mode estimation. To this aim, we posit that mode is not only constant throughout the musical piece, but also remains invariant by pitch transposition. Therefore, going back to the notations from Section 4: $T_{c}\bm{x}_{\mathrm{A}}$, $T_{c}\bm{x}_{\mathrm{B}}$, and $T_{c+k}\bm{x}_{\mathrm{B}}$ should elicit the same value of the mode estimator $\mu_{\bm{\theta}}$.
@@ 6.3 Results on key and mode classification
Training Semi-TONE on the full GSMK dataset yields a MIREX score of 72.6%; i.e., roughly on par with the supervised SOTA (73.1%). Figure 4 shows the confusion matrix of calibrated STONE and Semi-TONE on FMAK. Although our methods do not outperform the SOTA on key estimation, it brings insights into a novel framework that does not require high supervision for training. Moreover, we note that self-supervision remains beneficial even when the full GSMK dataset is available for training. Therefore, a promising avenue of research is to scale up the dataset of unlabeled recordings (see Section 4.1), thus widening the gap between Semi-TONE and Sup-TONE on FMAK.
@@ 7 Conclusion
STONE learns key signature profiles (KSP) via noncontrastive self-supervised learning in the time–frequency domain. We have seen than a semi-supervised extension of STONE (semi-TONE) reduces expert annotation by 90% less at no loss of MIREX score compared to the fully supervised variant (sup-TONE). The primary limitation of our work resides in the inability of the STONE objective (CPSD, i.e., cross-power spectral density) to distinguish major keys from minor keys.re work will study how STONE can be adapted to other pitch-relative MIR tasks.